OpenStudy (anonymous):

Some one please answer this. Its for a friend but only I have an account. Thanks :) Student tickets to the Homecoming game cost $5 each. General admission tickets cost $8 each. So far, 150 tickets have been sold. $900 has been collected. Write a system of equations for this model in standard form.

sammixboo (sammixboo):

\(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\) Now let's see if I can help :P

OpenStudy (anonymous):

whats the q?

OpenStudy (anonymous):

Student tickets to the Homecoming game cost $5 each. General admission tickets cost $8 each. So far, 150 tickets have been sold. $900 has been collected. Write a system of equations for this model in standard form.

OpenStudy (misty1212):

\[5x+8y=900\\ x+y=150\] should to it

OpenStudy (anonymous):

5x+8y=900 x+y=150